# Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option

### v4

@@ Abstract
This study examines strategic issues in fair rank-minimizing mechanisms, which choose an assignment that minimizes the average rank of object types to which agents are assigned and satisfy a fairness property called equal treatment of equals. As one of these fair mechanisms, the uniform rank-minimizing mechanism is considered. We focus on the case where agents can refuse their assignment and obtain the outside option instead. Without the refusal option, truth-telling is not strategically dominated by any strategies if a fair rank-minimizing mechanism is used. However, if agents have the option and the uniform rank-minimizing mechanism is used, then a strategy called an outside option demotion strategy strategically dominates truth-telling. Moreover, we show that adopting this strategy may lead to inefficient assignments. To counter this, we propose a modification of the uniform rank-minimizing mechanism, though it may lead agents to strategically reduce the number of acceptable types.
@@ 1 Introduction
In this study, we discuss the strategic issues of fair rank-minimizing allocation mechanisms mainly when agents have the option to refuse their assignment determined by one of the mechanisms and obtain their outside option instead. There are several agents and types of object, including a null object type, which corresponds to the outside option. In the mechanism, each agent reveals one preference ranking, which may or may not be their true preference ranking. In rank-minimizing mechanisms, assignments are determined as to minimize the total (or average) revealed rank of the objects to which the agents are assigned.
@@
The rank distributions of assignments are a critical consideration for many real-world matching authorities. For instance, Featherstone (2020) notes that Teach For America takes rank distribution into account when making assignment decisions. Consequently, rank-minimizing mechanisms are discussed in recent studies, including Featherstone (2020), Nikzad (2022), Ortega and Klein (2023), Feizi (2024), and Troyan (2024). Several of these studies highlight the advantages of rank-minimizing mechanisms, particularly regarding the efficiency of assignments. By the results of Featherstone (2020) and Feizi (2024), any rank-minimizing mechanism satisfies other efficiency properties, such as ordinal efficiency, which is first introduced by Bogomolnaia and Moulin (2001). Moreover, Ortega and Klein (2023) compare the outcomes of rank-minimizing mechanisms with those of the DA mechanism and the TTC mechanism, showing that rank-minimizing mechanisms improve upon the others in several aspects of efficiency.
@@
In this study, we especially consider the rank-minimizing mechanisms that satisfy a fairness property called equal treatment of equals, which requires that if two agents reveal the same preference ranking, then they are assigned to each object type with the same probability. This fairness property is discussed by Bogomolnaia and Moulin (2001) and many other previous studies on various allocation problems (Thomson 2011). We consider a slightly stronger version of this property. We call a rank-minimizing mechanism satisfying equal treatment of equals a fair rank-minimizing mechanism. To consider these mechanisms, we adopt the probabilistic assignment framework. Among the mechanisms, we first consider a representative one called the uniform rank-minimizing mechanism. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also mainly consider this mechanism, at least in the literature, this is the most common rank-minimizing mechanism.
@@
We discuss the strategic manipulation incentives of agents when the assignment is determined by a fair rank-minimizing mechanism. Specifically, we consider the case where agents can refuse their assignment, and if they refuse it, then the outside option is assigned to them. First, as a benchmark case, we discuss the situation where agents cannot refuse their assignment. This cas e is discussed by many previous studies, such as those mentioned above. Featherstone (2020) shows that any rank-minimizing mechanism is not strategy-proof. On the other hand, Troyan (2024) shows that the uniform rank-minimizing mechanism is not obviously manipulable, which is a strategic property introduced by Troyan and Morrill (2020). Moreover, Nikzad (2022) shows that the uniform rank-minimizing mechanism is Bayesian incentive compatible in uniform markets, which he primarily discusses. Related to but logically independent to their results, we show that, in this case, truth-telling is not strategically dominated by any untrue preference under any fair rank-minimizing mechanism. These results imply that in this case, strategic manipulation may not be a critical problem of rank-minimizing mechanisms.
@@
We consider an outside option demotion strategy, which involves revealing a ranking where the rank of the outside option is placed at the bottom, without changing the order of the truly acceptable object types. We show that, if the uniform rank-minimizing mechanism is used and agents have the refusal option, then any outside option demotion strategy weakly strategically dominates truth-telling. Moreover, with some additional conditions, some outside option demotion strategy (strictly) strategically dominates truth-telling.
The reason is as follows: first, for some preference profiles of other agents, an agent succeeds in avoiding the outside option and getting a truly acceptable object type because assigning the outside option to them significantly increases the sum of ranks. Of course, due to this manipulation, the agent faces the risk of being assigned an unacceptable object type. However, in the case where the agents have the refusal option, the agent can refuse any of such types and obtain the outside option instead. Moreover, we show that if an agent adopts an outside option demotion strategy and the uniform rank-minimizing mechanism is used, an inefficient assignment may be realized. This is because, in that case, an unacceptable type may be assigned to the agent who adopts the strategy, and then the agent refuses it, even though the type is desired by other agents. Therefore, in this case, strategic manipulation is a critical problem for rank-minimizing mechanisms.
@@ 2 Model
We consider the model discussed by Kojima and Manea (2010) and Featherstone (2020). Let $\mathcal{A}$ and $\mathcal{O}$ be finite sets of agents and object types respectively. We assume $\left|\mathcal{A}\right|\geq 2$ and $\left|\mathcal{O}\right|\geq 2$. Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$. Let $q_{o}\geq 1$ be the capacity of an object type $o\in\mathcal{O}$. We assume that there is a null object type denoted by $\varnothing\in\mathcal{O}$ satisfying $q_{\varnothing}\geq\left|\mathcal{A}\right|$. That is, the null object type can be assigned to all agents simultaneously. The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the first-best alternative object of an agent outside of this market. In Section 4, we consider the situation where the null object type has the other characteristic. To avoid trivial cases, we assume $q_{o}\in\left[1,\left|\mathcal{A}\right|\right)$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$. Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th-best one under $p_{a}$. For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$. Moreover, if $R\left(o,p_{a}\right)<R\left(\varnothing,p_{a}\right)$ (resp. $R\left(o,p_{a}\right)>R\left(\varnothing,p_{a}\right)$), then $o$ is said to be acceptable (resp. unacceptable) for $p_{a}$. Thus, $R\left(\varnothing,p_{a}\right)-1$ represents the number of acceptable object types for $p_{a}$. Let $r_{a}\in\mathcal{P}$ be the true preference order of agent $a$. Especially, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $o$ is said to be truly acceptable for $a$.
An assignment is represented as lotteries over types. Let $x_{ao}$ be the probability that agent $a$ assigns (a copy of) type $o$ object. Moreover, let $x$ be a matrix whose $\left(a,o\right)$ entry is equal to $x_{ao}$. We say that a matrix $x$ is an assignment if
@@
for all $a\in\mathcal{A}$ and all $o\in\mathcal{O}$. Let $\mathcal{X}$ be the set of all possible assignments. Specifically, let $y\in\mathcal{X}$ be a deterministic assignment if $y_{ao}$ is either 00 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$. Let $\mathcal{Y}\subseteq\mathcal{X}$ be a set of all possible deterministic assignments.
@@
We consider the preference of agent $a$ with $p_{a}$ with regard to two assignments $x$ and $x^{\prime}$. If
@@
Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$. For an assignment $x\in\mathcal{X}$ and a preference profile $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$, let $RV\left(x,p\right)$ be such that
@@
The efficiency property of ”rank-minimizing” is stronger than several efficiency properties, such as rank efficiency and ordinal efficiency, analyzed in this literature (see Figure 1 in Feizi (2024)). Among them, not being wasteful is a very weak property.
@@
Nikzad (2022) and Troyan (2024) consider a deterministic rank-minimizing assignment; that is, they focus on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$. However, by Corollary 1, such $y^{\ast}$ must also be rank-minimizing in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Moreover, this result allows us to show that the mechanism introduced later is well-defined.
@@
which is well-defined because $\sum\nolimits_{\bar{o}\in\mathcal{O}}q_{\bar{o}}>\left|\mathcal{A}\right|$. This means that for a given $p_{a}$, there are sufficient number of copies of object types that ranked higher than or equal to $\bar{k}\left(p_{a}\right)$. Trivially, if $p_{a}^{k}=\varnothing$, then $\bar{k}\left(p_{a}\right)\leq k$. We have the following result.
@@
Let $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$. Then, $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
Then, if an assignment is not wasteful, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$. Typically, in that case, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
@@ 3 Rank-minimizing Mechanism
In this section, we consider the following game. First, all agents simultaneously reveal a preference order in $\mathcal{P}$, which may or may not be their true preference order $r_{a}\in\mathcal{P}$. Let $p_{a}\in\mathcal{P}$ be the preference order revealed by $a\in\mathcal{A}$ and $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Second, based on $p$, an assignment is determined via a mechanism. We assume that all agents know the capacities of all object types $\left(q_{o}\right)_{o\in\mathcal{O}}$ and the mechanism used to determine the assignment, when they reveal a preference order. Note that in this section, the null object type is just that with sufficient number of copies.
A mechanism used in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$. First, we consider an efficiency property on mechanisms. We call $f$ a rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$.
Second, we introduce a fairness property. First, we introduce a well-known one. A mechanism $f$ is said to satisfy weak equal treatment of equals (hereafter weak ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}=p_{a^{\prime}}$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$. This requires that two agents revealing the same preference must have the same assignment in the ex-ante sense. This property is called just “equal treatment of equals” by many previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024). In this study, we mainly use a slightly stronger one. A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$ [^2] Han (2024) also considers a similar property, called symmetricity, which is slightly stronger than equal treatment of equals. Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects.
@@
By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th. Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
@@
We discuss strategic problems of fair rank-minimizing mechanisms. We call the strategy of revealing their true preference truth-telling. Featherstone (2020) shows that any rank-minimizing mechanism does not satisfy strategy-proofness. Therefore, in this paper, we consider a weaker strategic property.
We discuss whether truth-telling ($r_{a}$) can be strategically dominated by an untrue preference $p_{a}\neq r_{a}$. Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. By Lemma 2, if $k>\bar{k}\left(p_{a}\right)$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a fair rank-minimizing mechanism, because any result of a rank-minimizing mechanism is not wasteful. Thus, we focus only on the case where $k\leq\bar{k}\left(p_{a}\right)$ defined above; that is,
@@
By this lemma, we have the following result.
@@
Suppose that a fair rank-minimizing mechanism is used. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ such that revealing $p_{a}$ strategically dominates truth-telling.
This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference $p_{a}$, either $\left(f\left(r_{a},p_{-a}\right)\right)_{ao}=\left(f\left(p_{a},p_{-a}\right)\right)_{ao}$ for all $o\in\mathcal{O}$ and all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ or there must exist $p_{-a}$ such that $a$ is worse off by revealing it instead of $r_{a}$. Thus, if a fair rank-minimizing mechanism is used, any (essential) strategic manipulation by an agent carries some level of risk. This characteristic is quite weak, but in the subsequent section, we show that this is not satisfied when the agents can refuse their assignment and instead obtain the null object type.
Note that Proposition 2 holds for any fair rank-minimizing mechanisms but cannot be generalized to the rank-minimizing mechanisms satisfying weak ETE. We show this by introducing the following example.
@@
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$, $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$ and $q_{o_{1}}=1$. Let $\psi=\left(o_{1},o_{2},\varnothing,o_{3}\right)$ and $\psi^{\prime}=\left(o_{1},o_{2},o_{3},\varnothing\right)$. A rank-minimizing mechanism $f$ is such that
$\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi,\psi^{\prime},\psi\right)\right)_{a_{2}o_{1}}=\left(f\left(\psi,\psi,\psi^{\prime}\right)\right)_{a_{3}o_{1}}=0,$ (2)
$f\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$

### v5

@@ Abstract
This study examines strategic issues in fair rank-minimizing mechanisms, which choose an assignment that minimizes the average rank of object types to which agents are assigned and satisfy a fairness property called equal treatment of equals. As one of these fair mechanisms, the uniform rank-minimizing mechanism is considered. We focus on the case where agents can refuse their assignment and obtain the outside option instead. Without the refusal option, truth-telling is not strategically dominated by any strategies if a fair rank-minimizing mechanism is employed. However, if agents have the option and the uniform rank-minimizing mechanism is employed, then a strategy called an outside option demotion strategy strategically dominates truth-telling. Moreover, we show that adopting this strategy may lead to inefficient assignments. To counter this, we consider a modification of the uniform rank-minimizing mechanism, though it may lead agents to strategically reduce the number of acceptable types.
@@ 1 Introduction
In this study, we discuss the strategic issues of fair rank-minimizing allocation mechanisms mainly when agents have the option to refuse their assignment determined by one of the mechanisms and obtain their outside option instead. There are several agents and types of object, including a null object type, which corresponds to the outside option. In a mechanism, each agent reveals one preference ranking, which may or may not be their true preference ranking. In rank-minimizing mechanisms, assignments are determined as to minimize the total (or average) revealed rank of the objects to which the agents are assigned.
@@
The rank distribution of assignments is a critical consideration for many real-world matching authorities. For instance, Featherstone (2020) notes that Teach For America takes rank distribution into account when making assignment decisions. Consequently, rank-minimizing mechanisms are discussed in recent studies, including Featherstone (2020), Nikzad (2022), Ortega and Klein (2023), Feizi (2024), and Troyan (2024). Several of these studies highlight the advantages of rank-minimizing mechanisms, particularly regarding the efficiency of assignments. By the results of Featherstone (2020) and Feizi (2024), any rank-minimizing mechanism satisfies other efficiency properties, such as ordinal efficiency, which is first introduced by Bogomolnaia and Moulin (2001). Moreover, Ortega and Klein (2023) compare the outcomes of rank-minimizing mechanisms with those of the DA mechanism and the TTC mechanism, showing that rank-minimizing mechanisms improve upon the others in several aspects of efficiency.
@@
In this study, we focus on rank-minimizing mechanisms that satisfy the fairness property of equal treatment of equals, which ensures that agents revealing identical preference rankings are assigned to each object type with the same probability. [^1] Yokote et al. (2019) mention that this fairness property is a classical concept discussed by Aristotle in Book V of Nicomachean Ethics. This fairness property is discussed by Bogomolnaia and Moulin (2001) and many other previous studies on various allocation problems (Thomson 2011). We consider a slightly stronger version of this property. We call a rank-minimizing mechanism satisfying equal treatment of equals a fair rank-minimizing mechanism. To consider these mechanisms, we adopt the probabilistic assignment framework. Among the mechanisms, we first consider a representative one called the uniform rank-minimizing mechanism. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also mainly consider this mechanism, at least in the literature, this is the most common rank-minimizing mechanism.
@@
We discuss the strategic manipulation incentives of agents when the assignment is determined by a fair rank-minimizing mechanism. Specifically, we consider the case where agents can refuse their assignment, and if they refuse it, then the outside option is assigned to them. First, as a benchmark case, we discuss the situation where agents cannot refuse their assignment. This case is discussed by many previous studies, such as those mentioned above. Featherstone (2020) shows that any rank-minimizing mechanism is not strategy-proof. On the other hand, Troyan (2024) shows that the uniform rank-minimizing mechanism is not obviously manipulable, which is a strategic property introduced by Troyan and Morrill (2020). Moreover, Nikzad (2022) shows that the uniform rank-minimizing mechanism is Bayesian incentive compatible in uniform markets, which he primarily discusses. Related to but logically independent to their results, we show that, in this case, truth-telling is not strategically dominated by any untrue preference under any fair rank-minimizing mechanism. These results imply that in this case, strategic manipulation is a critical problem of rank-minimizing mechanisms.
@@
We consider an outside option demotion strategy, which involves revealing a ranking where the rank of the outside option is placed at the bottom, without changing the order of the truly acceptable object types. We show that, if the uniform rank-minimizing mechanism is employed and agents have the refusal option, then any outside option demotion strategy weakly strategically dominates truth-telling. Moreover, with some additional conditions, some outside option demotion strategy (strictly) strategically dominates truth-telling.
The reason is as follows: First, for some preference profiles of other agents, an agent can successfully avoid the outside option and obtain a truly acceptable object type because assigning them the outside option would significantly increase the sum of ranks. However, this manipulation carries the risk of being assigned an unacceptable object type. Yet, when agents have the refusal option, they can simply reject such assignments and choose the outside option instead. Thus, in this case, the risk associated with adopting this strategy is entirely eliminated.
Moreover, we show that if an agent adopts an outside option demotion strategy and the uniform rank-minimizing mechanism is employed, an inefficient assignment may be realized. This is because, in that case, an unacceptable type may be assigned to the agent who adopts the strategy, and then the agent refuses it, even though the type is desired by other agents. Therefore, in this case, strategic manipulation is a critical problem for rank-minimizing mechanisms.
@@ 2 Model
We consider the model discussed by Kojima and Manea (2010) and Featherstone (2020). Let $\mathcal{A}$ and $\mathcal{O}$ be finite sets of agents and object types respectively. We assume $\left|\mathcal{A}\right|\geq 2$ and $\left|\mathcal{O}\right|\geq 3$. Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$. Let $q_{o}\geq 1$ be the number of copies (the capacity) of an object type $o\in\mathcal{O}$. We assume that there is a null object type denoted by $\varnothing\in\mathcal{O}$ satisfying $q_{\varnothing}\geq\left|\mathcal{A}\right|$. That is, the null object type can be assigned to all agents simultaneously. The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the first-best alternative object of an agent outside of this market. In Section 4, we consider the situation where the null object type has the other characteristic. To avoid trivial cases, we assume $q_{o}\in\left[1,\left|\mathcal{A}\right|\right)$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
Let $\mathcal{P}$ be the set of all possible strict rankings of the elements in $\mathcal{O}$. Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th-best one under $p_{a}$. For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$. Moreover, if $R\left(o,p_{a}\right)<R\left(\varnothing,p_{a}\right)$ (resp. $R\left(o,p_{a}\right)>R\left(\varnothing,p_{a}\right)$), then $o$ is said to be acceptable (resp. unacceptable) for $p_{a}$. Thus, $R\left(\varnothing,p_{a}\right)-1$ represents the number of acceptable object types for $p_{a}$. Let $r_{a}\in\mathcal{P}$ be the true preference order of agent $a$. Especially, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $o$ is said to be truly acceptable for $a$.
An assignment is represented as lotteries over types. Let $x_{ao}$ represent the probability that agent $a$ assigns (a copy of) type $o$ object. Moreover, let $x$ be a matrix whose $\left(a,o\right)$ entry is equal to $x_{ao}$. We say that a matrix $x$ is an assignment if
@@
for all $a\in\mathcal{A}$ and all $o\in\mathcal{O}$. Let $\mathcal{X}$ be the set of all possible assignments. Specifically, $y\in\mathcal{X}$ is said to be a deterministic assignment if $y_{ao}$ is either 00 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$. Let $\mathcal{Y}\subseteq\mathcal{X}$ be a set of all possible deterministic assignments.
@@
We consider the preference of agent $a$ denoted by $p_{a}$ with regard to two assignments $x$ and $x^{\prime}$. If
@@
Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$. In words, an assignment is considered wasteful for a given preference profile if there is at least one agent who, despite the fact that a more desirable object is still available, does not receive as much of it as they could, and instead receives some amount of a less preferred object.
For an assignment $x\in\mathcal{X}$ and a preference profile $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$, let $RV\left(x,p\right)$ be such that
@@
The efficiency property of rank-minimizing is stronger than several other efficiency properties, such as rank efficiency and ordinal efficiency, analyzed in this literature (see Figure 1 in Feizi (2024)). Among them, not being wasteful is a very weak property.
@@
Nikzad (2022) and Troyan (2024) focus only on deterministic rank-minimizing assignments; that is, they consider a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$. However, by Corollary 1, such $y^{\ast}$ must also be rank-minimizing in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Moreover, this result allows us to show that the mechanism introduced later is well-defined.
@@
which is well-defined because $\sum\nolimits_{\bar{o}\in\mathcal{O}}q_{\bar{o}}>\left|\mathcal{A}\right|$. This means that for a given $p_{a}\in\mathcal{P}$, there are sufficient number of copies of object types that ranked higher than or equal to $\bar{k}\left(p_{a}\right)$. Trivially, if $p_{a}^{k}=\varnothing$, then $\bar{k}\left(p_{a}\right)\leq k$. We have the following result.
@@
If $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$, then $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
This means that if an assignment is not wasteful for $p$, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$. Typically, under a nonwasteful assignment, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
@@ 3 Rank-minimizing Mechanism
In this section, we consider the following game. First, all agents simultaneously reveal a preference order in $\mathcal{P}$, which may or may not be their true preference order $r_{a}\in\mathcal{P}$. Let $p_{a}\in\mathcal{P}$ be the preference order revealed by $a\in\mathcal{A}$ and $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Second, based on $p$, an assignment is determined via a mechanism.
We assume that all agents know the capacities of all object types $\left(q_{o}\right)_{o\in\mathcal{O}}$ and the mechanism is employed to determine the assignment, when they reveal a preference order. Note that in this section, the null object type is just that with sufficient number of copies.
A mechanism employed in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$. First, we consider an efficiency property on mechanisms. We call $f$ a rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$.
Second, we introduce a fairness property. First, we introduce a well-known one. A mechanism $f$ is said to satisfy weak equal treatment of equals (hereafter weak ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}=p_{a^{\prime}}$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$. This requires that two agents revealing the same preference must have the same assignment in the ex-ante sense. This property is called just “equal treatment of equals” by many previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024). In this study, we mainly use a slightly stronger one.
A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$ [^3] Han (2024) also considers a similar property, called symmetricity, which is slightly stronger than equal treatment of equals. Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects. By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th. Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
@@
We discuss strategic problems of fair rank-minimizing mechanisms. We call the strategy of revealing their true preference truth-telling. Featherstone (2020, Proposition 10) shows that any rank-minimizing mechanism does not satisfy strategy-proofness. Therefore, in this paper, we consider a weaker strategic property.
For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly strategically dominated by (revealing) $p_{a}$ if $a$ weakly prefers $f\left(p_{a},p_{-a}\right)$ to $f\left(p_{a}^{\prime},p_{-a}\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) strategically dominated by (revealing) $p_{a}$ if $p_{a}^{\prime}$ is weakly strategically dominated by $p_{a}$ and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(p_{a},p_{-a}^{\prime}\right)$ to $f\left(p_{a}^{\prime},p_{-a}^{\prime}\right)$.
We discuss whether truth-telling ($r_{a}$) can be strategically dominated by revealing an untrue preference $p_{a}\neq r_{a}$. Fix any $r_{a}$ and any $p_{a}\neq r_{a}$. Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. By Lemma 2, if $k>\bar{k}\left(p_{a}\right)$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a fair rank-minimizing mechanism, because any result of a rank-minimizing mechanism is not wasteful. Thus, we focus only on $r_{a}$ and $p_{a}$ satisfying $k\leq\bar{k}\left(p_{a}\right)$ defined above; that is,
@@
By Lemma 3, we have the following result.
@@
Suppose that a fair rank-minimizing mechanism is employed. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ such that revealing $p_{a}$ strategically dominates truth-telling.
This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference $p_{a}$, either $\left(f\left(r_{a},p_{-a}\right)\right)_{ao}=\left(f\left(p_{a},p_{-a}\right)\right)_{ao}$ for all $o\in\mathcal{O}$ and all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ or there must exist $p_{-a}$ such that $a$ is worse off by revealing it instead of $r_{a}$. Thus, if a fair rank-minimizing mechanism is employed, any (essential) strategic manipulation by an agent carries some level of risk. This characteristic is quite weak, but in the subsequent section, we show that this is not satisfied when the agents can refuse their assignment and instead obtain the null object type.
Note that Proposition 2 holds for any fair rank-minimizing mechanisms but cannot be generalized to rank-minimizing mechanisms satisfying weak ETE. We show this by introducing the following example.
@@
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$, $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$ and $q_{o_{1}}=1$. Let $\alpha=\left(o_{1},o_{2},\varnothing,o_{3}\right)$ and $\alpha^{\prime}=\left(o_{1},o_{2},o_{3},\varnothing\right)$. Let a rank-minimizing mechanism $f^{\ast}$ be such that, if $\left(p_{a_{i}},p_{a_{j}},p_{a_{k}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$ where $i,j,k=1,2,3$ and they are distinct, then
$\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{i}o_{1}}$ $=$ $0,$ (2)
$f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$